OpenStudy (maddy1251):

g(x)= x^3-2x^2+x/x^2-1 Horizontal asymptote, oblique asymptote, hole, domain, range is what I need to find I am having trouble. Will fan and medal.

OpenStudy (superdavesuper):

u can try typing the function in the Wolfram web site...

OpenStudy (superdavesuper):

it should be like this right: g(x)= (x^3-2x^2+x)/(x^2-1)

OpenStudy (maddy1251):

There it is

OpenStudy (superdavesuper):

hahahaaa u trying 2 confused me? @Maddy1251 :P

OpenStudy (maddy1251):

No :p

OpenStudy (chillout):

Do you know how to find oblique, horizontal and vertical asymptotes?

OpenStudy (anonymous):

factor and cancel as a first step

OpenStudy (chillout):

For the vertical asymptotes, set the denominator to 0 (x²-1). For the oblique asymptote, effectuate long division until you find the equation of a line.

OpenStudy (anonymous):

\[\frac{x(x-1)(x-1)}{(x+1)(x-1)}\] then cancel the common factor of \(x-1\) top and bottom

OpenStudy (anonymous):

since you are cancelling the \(x-1\) term there is "hole" at \(x=1\)

OpenStudy (anonymous):

there is no vertical asymptote however

OpenStudy (anonymous):

once you get to \[\frac{x(x-1)}{x+1}\] you find the "oblique" asymptote by division you get a quotient and a remainder when you divide ignore the remainder, the quotient is the "oblique" asymptote

OpenStudy (anonymous):

there is also a vertical asymptote set \(x+1=0\) get \(x=-1\) the vertical line \(x=-1\) is the vertical asymptote

OpenStudy (chillout):

@satellite73, rewriting it wouldn't work only for calculating limits? Are you sure there is not an asymptote at x=-1 and an oblique asymptote y=x-2?

OpenStudy (anonymous):

ok i lied, there is a vertical asymptote at \(x=-1\) there is no horizontal one

OpenStudy (maddy1251):

@satellite73 I thought the hole was at -1?

OpenStudy (maddy1251):

@ChillOut I thought the same

OpenStudy (chillout):

@Maddy1251, rewriting it like that only works for calculating limits. x=-1 is a vertical asymptote.

OpenStudy (chillout):

There are no holes whatsoever in this function. But there is an oblique asymptote (which in turn makes horizontal asymptotes "impossible" to exist). Just remember that whenever the numerator is one degree higher than the denominator there will always be a slant asymptote!

OpenStudy (maddy1251):

@ChillOut So, the hole was found how exactly?

OpenStudy (maddy1251):

@ChillOut the cubed x makes it slant because the x squared on the bottom, yes?

OpenStudy (maddy1251):

And how isn't there a hole for the graph?

OpenStudy (chillout):

Generally speaking, holes exist when you get 0/0.

OpenStudy (chillout):

@Maddy1251, yes. You have a cubic function on top of a squared one. So, yeah, you get a slant asymptote.

OpenStudy (maddy1251):

x=-1 would produce that.

OpenStudy (maddy1251):

Gotcha, that makes sense.

OpenStudy (chillout):

Remember when you rewrite a function to calculate a limit and that limit exists? That's how you find a hole.

OpenStudy (chillout):

Let's say, f(x)=\(\begin{cases} x+1\:\: if\:\: x≠ \: 5 \\ 10\:\: if\:\:x=5 \end{cases}\) There is a hole at x=5

OpenStudy (maddy1251):

@ChillOut I just did the math and you are right, my bad terribly. Lol. 1 cubed = -1, then its -1-2=-3 then -3+(-1)=-2/ then the bottom would equal 0, soooo -2/0 is undefined, so yeah you're right. That's how I checked. The limited function I kinda know. I see.

OpenStudy (chillout):

Any questions?

OpenStudy (maddy1251):

So the horizontal does exist because of the oblique?

OpenStudy (chillout):

Doesn't*.

OpenStudy (maddy1251):

Thats what I meant. I can only type with one hand, I was in a riding accident lol, but okay. I think I may have this. Satelite factoring really helped actually, but you are very smart in this. Your help was really really useful

OpenStudy (chillout):

No problem, :P. Glad to help!

OpenStudy (maddy1251):

Sorry I couldnt medal, @superdavesuper got you I think. I fanned @ChillOut

OpenStudy (chillout):

I'm not here for medals. Do not worry about that :).

OpenStudy (maddy1251):

You're awesome!! Thank youu :)